How do u put x+3y=3 into function for
Answer by stanbon(75887)

How do u put x+3y=3 into function form?
------------------
x + 3y = 3
-------
Solve for "y"::
3y = -x + 3
------
y = (-x+3)/3
-----
y = (-1/3)x + 1
------------------
Cheers,
Stan H.
----------